A Discussion of the magical attributes of the hero in Fenian Literature with specific reference to the tale The Pursuit of Diarmuid and Gráinne

GA4115 Essay

Oliver Gerler 9954465 GA4115: Irish Language 1 University of Limerick, Ireland December 14, 1999

What is magic? How do we know what magic is like? How does magic define itself? How do we know when magic occurs? How does magic know how to occur that we can percept it?

As Arthur C. Clarke puts it ”Any sufficiently advanced technology is indistinguishable from magic.” it will become very relative. Magic soon loses its absolute value. So what is magic then? Does it perhaps mean something different for each of us? Certainly it does, I would say, as none of us has the same level of knowledge as one other has.

In this sense magic changes relatively to our knowledge of how to do certain things, or how much we know about certain facts that are not that widespread. It is just because people do not understand. But do they want to understand? They just want to get enchanted, they want to be able to be astonished like children or Alice in Wonderland.

Ever since the Fenian Cycle has passed by, the stories and tales from the old times have fascinated the people, and still do. The longer the time has gone the more imposing and heroic the tales will grow. A mixture of different times and epochs will happen, though, but some can argue if this diminishes the effect the tale has to the listeners if it is told properly or read in the Irish original or a good translation.

How does this relate to the events taking place in The Pursuit of Diarmuid and Gráinne (Tóruigheacht Dhiarmada agus Ghráinne)? The Fenian Cycle describes the island of Ireland of around 300 A.D. in tales mostly written from 1200 onwards, usually in form of ballads or lays. Despite the saga of the warriors of the Fianna^{1}, the warrior band around Fionn son of Cumhaill, has been christianized, many elements from earlier, mainly the mythological, cylces can be found. This can for example be seen in the conversations between (the at least 300 years old) Oisín and St.Patrick^{2} as well as in the occurrence of Aonghus an Brogha, Aengus of the Brugh, one of the Tuatha de Danaan and Diarmuid’s foster-father. The characters in the tales are depicted as being very tall and strong and hence live a lot longer, as also Caoilte seems to be around at the time of St.Patrick.

As the usual lifespan of people at that time cannot have been 300 years this for sure has to be treated as a magic element of the tales of the cycle. The appearances of Aonghus an Brogha, of whom is said that ”he proceeded in the company of pure cold wind […], and he went, unknown to Fionn and to the Fiana of Ireland […]”^{3}, have to be counted to magic, too.

Diarmuid ó Duibhne, as his full name is, is said to have had a love-spot on his face or his forehead^{4}. This love-spot forces women to fall in love with him at the first sight, they get attracted to him magically. This love-spot causes the start of things going amiss in The Pursuit of Diarmuid and Gráinne: in the literary form just because it is there and Diarmuid has the effect on women that they fall in love with him and act irrationally on that. In the oral tradition it is said, that Gráinne accidentally sees the love-spot and therefore starts to let the sleep-potion go round.

1It is said, that ”Their strength was seven-score and ten officers, each man of these having thrice nine warriors […]”. See [7, p. 95].

2The Dialogue between Oisín and Patrick [4, pp. 204ff.].

3See [6, lines 361-367, lines 379-380 and several others].

4See [2, p. 228].

The magic in this is that people are not able to understand how or even why the love-spot and the sleep-potion is effective. But understanding is not required. If someone learns to understand how magic works, this will no longer be magic anymore. It all relates to the relative level of knowledge oneself gained in his life.

The story continues by Gráinne putting a geis, an injunction or taboo, on Diarmuid to elope with him and leave his band, his friends and Fionn behind. These geasa are very powerful weapons, they can convert proud men to obedient and submissive followers, who do not question anything. Throughout the whole story of The Pursuit of Diarmuid and Gráinne it can be seen that Gráinne just needs to bend the proverbial finger and Diarmuid jumps up and fulfills her wishes.

But geasa cannot only be cast, they can happen by curse and even without knowledge of the person who falls under the spell^{5}, which very likely can lead this person in unsolveable problems, as in The Pursuit of Diarmuid and Gráinne where in the end it leads to the death of Diarmuid.

The question is if all these injunctions are of magic nature? The one that Gráinne puts on Diarmuid quite sure is, because he has to go with her against his own will^{6}. As it comes to the end of the story Diarmuid is told by Fionn on the mound where they meet to face the boar that it is one of the prohibitions put on Diarmuid (by his foster-father) to hunt pigs. As there is another prohibition that Diarmuid has to have the same length of life as this pig, Aonghus forbade him to hunt swine at all. Yet, Diarmuid stays on top of the mound and both, he and the boar, get killed in the end.

The problems a character in the stories has to face occur from contradicting spells cast on him. Gráinne and Diarmuid are not able to get out by the main doors after the banquet as Fionn has the keys, but Gráinne knows of a backdoor. yet, Diarmuid is not able to go through this door, as it is another geis on him, not to go through any escape-door whatsoever. It is only because the heroes and battle-champions are not only said, but truly able to go in and out of a fort by leaping over the palisades ”by the shafts of his javelins and the staves of his spears”^{7}.

But what would have happened, if Diarmuid could not escape through another magic ”function” of him and neither by the escape-door? He has to follow the geis that they elope together. The question arises what the injunctions and geasa are for? Is it only to bring some interesting elements into the story that the reader will stick to it?

The story as such can only happen because of the effect and consequences the spelling of a geis had. Not even a magically ”improved” hero could ignore their impact, because of their magic nature. On the other hand, the concept of honour and face had meaning and huge value in the society. There was no easy way to by-pass the rituals that were laid on one if that would have meant conflict with codices on how to behave within the society. Therefore not all of the injections, prohibitions and taboos had to be magical to have an effect as if they would have been so.

5See [6, lines 1447ff.].

6It has to be noted that there are version of The Pursuit of Diarmuid and Gráinne which do not mention the love-spot or the injunction at all. See [6, p. xxvii].

7See [6, lines 198-202].

The geasa are not the only magic attributes in the stories. A lot more are present, many of them only existing because of the tallness and immense strength of the heroes, though. Many stones and rocks are said to have been thrown by the giants and heroes of the older days. The giant dolmens all around the country are said to be the ”leaba Dhiarmada agus Ghráinne”. But one particular speciality of Diarmuid is his abilty to do wide and long jumps. Most of the times in The Pursuit of Diarmuid and Gráinne he only escapes because he leapes over the gathered Fianna and finds himself outside the ring of men which should have kept him held.

But Diarmuid is not the only one who has magic abilities. He is one out of many. Fionn had his magic thumb of knowledge^{8} and he had prophetic abilities^{9}. Oisín, Fionn’s son, had a close connection to otherworld^{10}. Conán had the ability to kill hundreds (but he did not know how to do it).

The existence of the hero itself has to be treated as magic as they are said not only to be extraordinarily tall but also have a much bigger life-span. The heroes have to be seen as supernatural personages and not as mere mortals.

Animals have been regarded as having magic abilities, too. So for example Fionn’s hound Bran, who could understand the language of the people^{11}, or the story of the magic pig^{12}. Interestingly enough Cork’s oral tradition claims that Fionn had a dozen cats^{13}.

Even dead things like cloths should have had some magic abilities. There is the story of the magic cloak^{14} a woman brings with her as she visits the Fianna. It tests the modesty of the woman who puts it on, but as all but one women of the Fianna fail the test, Fionn bids the woman of the magic cloak begone. It can be argued here, that it was the woman who actually did the magic not the cloak.

As it comes to the end of the story of The Pursuit of Diarmuid and Gráinne it is not entirely clear in which way it ended in its original version. Much material covering the involvements of the Fianna is still preserved^{15} and the stories are retold in newer times, not always in quite the same heroic and epic manner^{16}. On the other hand there are very accurate translations of the Irish Texts Society^{17} which are successful in their attempt to convert the Irish stories into the English language and still keep the feeling of epos and heroism alive.

8Which in [6, lines 1499-1503] has become the tooth of knowledge, where Fionn says, he had to ”[…] put my thumb under my tooth of knowledge, and knowledge and truth were revealed to me […]”.

9Fionn’s Prophecy [4, pp. 154ff.] and others.

10See ”Oisin in Tír na nOg” (Oisín in the land of the forever young).

11See [6, lines 297-300]

12The Magic Pig [4, pp. 184ff.].

13”slender cat, old cat, fierce cat, excellent cat, fat cat, morbid cat, striped cat, black cat, grey cat, wild cat, yearling cat and malingerer.” [2, p. 220].

14The Magic Cloak [4, pp. 330 ff.].

15An enumeration of the many manuscripts of The Pursuit of Diarmuid and Gráinne alone can be found in [6, pp. xiv f.].

In comparing these manuscripts^{18} it can be seen that there is no doubt on the way Diarmuid dies but what follows is subject to discussion^{19}. In most of the versions Gráinne sends for her children. As this will have had a purpose it seems plausible that it is to exhort them to avenge their father’s death.

This open ending provides for some continuation, as for example is done in the Early Modern poem on the daughter of Diarmaid, Éachtach^{20}.

Several other endings and possible motifs for sequels come to mind and the different existing endings themselves help to keep the interest and curiosity maintained. So the magic continues and the reader is still enchanted in the world of tales, stories and myths as it should be.

16See [3, pp. 180ff.] for a quite unexciting version of The Pursuit of Diarmuid and Gráinne.

17See [6].

18See [6, p. xiv ff.].

19According to [6, p. xvii] there are five different endings. One of them is of particular interest as it tells of Gráinne marrying Fionn after the death of Diarmuid.

20It tells of the fierce war between Diarmuid’s children and Fionn with the result that Éachtach was killed and Fionn badly injured.

Bibliography

[1] Gerler Oliver, Lecture notes to GA4115, University of Limerick, 1999

[2] Ó hÓgáin Dáithí, Magic attributes of the hero in Fenian lore, Bealoidas vols. 54/55, Folklore of Ireland Society, 1986/87, ISSN 0332-270X

[3] Heaney Marie, Over Nine Waves - A Book of Irish Legends, faber and faber Ltd., 1995, ISBN 0-571-17518-X

[4] Murphy Gerard, ed., Duanaire Finn - The Book of the Lays of Fionn, Part II, Irish Texts Society, vol. XXVIII, 1933

[5] Murphy Gerard, ed., Duanaire Finn - The Book of the Lays of Fionn, Part III, Irish Texts Society, vol. XLIII, 1953

[6] Ní Shéaghdha Nessa, ed., Tóruigheacht Dhiarmada agus Ghráinne - The Pursuit of Diarmaid and Gráinne, Irish Texts Society, vol. XLVII, 1967

[7] Greene David H., ed., An Anthology of Irish Literature, vol.I, New York University Press, 1954, repr. 1971, ISBN 0-8147-2953-3